The first day of class, you saw an apparatus with two syringes having different diameters connected by a tube. You observed that to maintain equilibrium you had to push harder on the tube with the larger diameter than on the tube with the smaller diameter. Explain why that happened.
Answer #1

According to Bernoulli's Theorem in Fluid mechanics

P1 = P2 means for the same liquid in a pipe pressure at two different points will be the same.

Means,

F1/A1 = F2/A2   :: F1/F2 = A1/A2

Let us assume A1 > A2

Then A1/A2 > 1

Means F1/F2 >1

Means F1 > F2

Means we have to apply larger force at larger diameter portion to maintain the equillibrium